# Consistency Models for Scalable and Fast Simulation-Based Inference

### v2

@@ Abstract
Simulation-based inference (SBI) is constantly in search of more expressive algorithms for accurately inferring the parameters of complex models from noisy data. We present consistency models for neural posterior estimation (CMPE), a new free-form conditional sampler for scalable, fast, and amortized SBI with generative neural networks. CMPE combines the advantages of normalizing flows and flow matching methods into a single generative architecture: It essentially distills a continuous probability flow and enables rapid few-shot inference with an unconstrained architecture that can be tailored to the structure of the estimation problem. Our empirical evaluation demonstrates that CMPE not only outperforms current state-of-the-art algorithms on three hard low-dimensional problems, but also achieves competitive performance in a high-dimensional Bayesian denoising experiment and in estimating a computationally demanding multi-scale model of tumor spheroid growth.
Simulation-Based Inference, Bayesian Inference, Probabilistic Machine Learning, Consistency Models
[…]
@@ 1 Introduction
Computer simulations play a fundamental role across countless scientific disciplines, ranging from physics to biology, and from climate science to economics (Lavin et al., 2021). Simulation programs can generate observables $\mathbf{x}\in\mathcal{X}$ as a function of unknown parameters $\boldsymbol{\theta}\in\Theta$ and latent program states $\boldsymbol{\xi}\in\Xi$ (Cranmer et al., 2020). This forward problem is typically well-understood through scientific theories instantiated as generative models.
The inverse problem, however, is much harder, and forms the crux of Bayesian (probabilistic) inference: Reasoning about the unknowns $\boldsymbol{\theta}$ based on observables $\mathbf{x}$. Bayes’ theorem captures the full posterior distribution of plausible parameters $\boldsymbol{\theta}$ conditional on the observed data $\mathbf{x}$ as $p(\boldsymbol{\theta}\,|\,\mathbf{x})\propto p(\boldsymbol{\theta})\,p(\mathbf{x}\,|\,\boldsymbol{\theta})$, given a prior $p(\boldsymbol{\theta})$. While generating synthetic data from the simulation program is possible (albeit potentially slow), the likelihood density $p(\mathbf{x}\,|\,\boldsymbol{\theta})$ is generally not explicitly available. Instead, it is only implicitly defined through a high-dimensional integral over all possible program execution paths (Cranmer et al., 2020). A new era of simulation intelligence (Lavin et al., 2021) addresses this challenging setting with remarkable success across application domains (Butter et al., 2022; Radev et al., 2021b; Shiono, 2021; Bieringer et al., 2021; von Krause et al., 2022; Ghaderi-Kangavari et al., 2023).
These advances in simulation-based inference (SBI) are co-developing with rapid progress in conditional generative modeling.
Figure 1: Inverting a probability flow on a target vector $\boldsymbol{\theta}$ forms the core of modern denoising models. Consistency models (Song et al., 2023) learn a shortcut $f_{\phi}$ that maps any point on the trajectory directly to the origin $\boldsymbol{\theta}_{0}$. We extend consistency models by conditioning on observable data $\mathbf{x}$ and infer mechanistic model parameters $\boldsymbol{\theta}$ as the target vector. As a consequence, the probability distribution over the origin $p({\color[rgb]{1,0.5234375,0.5234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.5234375,0.5234375}\pgfsys@color@rgb@stroke{1}{0.5234375}{0.5234375}\pgfsys@color@rgb@fill{1}{0.5234375}{0.5234375}\boldsymbol{\theta}_{0}}\,|\,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{x}})$ is the Bayesian posterior by definition. We call this approach consistency model posterior estimation (CMPE), which is a new amortized method for scalable and fast simulation-based inference.
Currently, advancements in generative neural networks primarily manifest in remarkable performance on generating images (Goodfellow et al., 2014), videos (Vondrick et al., 2016), and audio (Engel et al., 2020).
@@
Within the class of generative models, score-based diffusion models (Song et al., 2021; Rombach et al., 2022; Ho et al., 2020; Batzolis et al., 2021) have recently attracted significant attention due to their sublime performance as realistic generators. Diffusion models are strikingly flexible but require a multi-step sampling phase to denoise samples (Song et al., 2021). To address this shortcoming, Song et al. (2023) proposed consistency models, which are trained to perform few-step generation by design. In this paper, we port consistency models to simulation-based inference (see Figure 1) to achieve an unprecedented combination of both scalable and fast neural posterior estimation. Our contributions are:
@@
    We illustrate the fundamental advantages of consistency models for simulation-based inference: expressive free-form architectures and fast inference;
@@
    We demonstrate that CMPE outperforms both normalizing flows and flow matching on three benchmark experiments (see Figure 2 above), high-dimensional Bayesian denoising, and a tumor spheroid model.
@@ 2 Preliminaries and Related Work
This section recaps simulation-based inference via normalizing flows, flow matching, and score-based diffusion models.
@@ 2.1 Simulation-Based Inference (SBI)
The defining property of SBI methods is that they rely only on the ability to sample from the data-generating process, $p(\mathbf{x},\boldsymbol{\theta})$, as opposed to likelihood-based methods which rely on the ability to evaluate the likelihood $p(\mathbf{x}\,|\,\boldsymbol{\theta})$. Closely related, probabilistic neural methods for SBI can be classified into sequential (Papamakarios & Murray, 2016; Lueckmann et al., 2017; Greenberg et al., 2019; Durkan et al., 2020; Deistler et al., 2022; Sharrock et al., 2022) vs. amortized (Ardizzone et al., 2019; Gonçalves et al., 2020; Radev et al., 2020; Pacchiardi & Dutta, 2022; Avecilla et al., 2022; Dax et al., 2023b; Geffner et al., 2022; Radev et al., 2021a; Elsemüller et al., 2023b) methods.
Sequential inference algorithms iteratively refine the prior $p(\boldsymbol{\theta})$ to generate simulations in the vicinity of the target observation. Thus, they are not amortized, as each new observed data set requires a costly re-training of the neural approximator tailored to the particular target observation. In contrast, amortized methods train the neural approximator to generalize over the entire prior predictive space of the model. This allows us to query the approximator with any new data set that is presumed to originate from the model’s scope. In fact, amortization can be performed across any component of the model, including multiple data sets (Gonçalves et al., 2020) and contextual factors, such as the number of observations in a data set (Radev et al., 2020), heterogeneous data sources (Schmitt et al., 2023b), or even different probabilistic models and data configurations (Elsemüller et al., 2023a). Our current work is situated in the amortized setting.
@@ 2.2 Normalizing Flows for Neural Posterior Estimation
Neural posterior estimation (NPE) methods for SBI have traditionally relied on conditional normalizing flows for learning a neural density estimator $p_{\boldsymbol{\phi}}(\boldsymbol{\theta}\,|\,\mathbf{x})$ from simulated pairs $(\boldsymbol{\theta},\mathbf{x})$ of parameters and data (Rezende & Mohamed, 2015; Papamakarios et al., 2021). The learnable neural network weights are subsumed as $\boldsymbol{\phi}$. A normalizing flow defines a single-shot transformation between the parameters $\boldsymbol{\theta}$ and a latent variable $\mathbf{z}$ with a simple base distribution (e.g., Gaussian). Normalizing flows can learn closed-form approximate densities via maximum likelihood training:
@@
Two prominent normalizing flow architectures in SBI are affine coupling flows (ACF; Dinh et al., 2016) and neural spline flows (NSF; Durkan et al., 2019). A major disadvantage of these architectures is the invertibility requirement: It restricts the design space of the neural networks to invertible functions with cheap Jacobian calculations. In the following, we will briefly recap recent works that have pioneered the use of multi-step, free-form architectures for SBI, inspired by the success of such models in a variety of generative tasks (Batzolis et al., 2021; Rombach et al., 2022).
Table 1: Comparison of neural posterior estimation methods, adapted from Dax et al. (2023b). NPE: neural posterior estimation; NPSE: neural posterior score estimation; FMPE: flow matching posterior estimation; CMPE: consistency model posterior estimation (Ours).
NPE NPSE FMPE CMPE (Ours)
[…]
Unconstrained networks No Yes Yes Yes
Network passes for sampling Single Many Many Few
${}^{1}$ee Section 3.4 for details.
@@ 2.3 Flow Matching for Posterior Estimation
Dax et al. (2023b) applied flow matching techniques (Lipman et al., 2023; Liu et al., 2022) in SBI, an approach called flow matching posterior estimation (FMPE). Flow matching lifts the invertibility restrictions and enables the use of free-form neural networks. FMPE is based on optimal transport, where the mapping between base and target distribution is parameterized by a continuous process driven by a vector field on the sample space for each time step $t\in[0,1]$. Here, the distribution at $t=1$ is a unit Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I})$ and the distribution at $t=0$ is the target posterior. The FMPE loss replaces the maximum likelihood term in the NPE objective (Eq. 1) with a conditional flow matching objective,
@@
where $\mu_{\boldsymbol{\phi}}$ denotes the conditional vector field parameterized by a free-form neural network with trainable parameters $\boldsymbol{\phi}$, and $u_{t}$ denotes a marginal vector field, which can be as simple as $u_{t}(\boldsymbol{\theta}_{t}\,|\,\boldsymbol{\theta})=\boldsymbol{\theta}_{1}-\boldsymbol{\theta}$ for all $t$ (Liu et al., 2022). We obtain posterior draws $\boldsymbol{\theta}_{0}\sim p(\boldsymbol{\theta}\,|\,\mathbf{x})$ by solving $\mathrm{d}\boldsymbol{\theta}_{t}=-\mu(\boldsymbol{\theta}_{t},t;\mathbf{x})$ in reverse on $t\in[0,1]$, starting with noise samples $\boldsymbol{\theta}_{1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. We can use any off-the-shelf ODE solver for transforming noise $\boldsymbol{\theta}_{1}$ into a draw $\boldsymbol{\theta}_{0}$ from the approximate posterior. In principle, the number of steps $K$ in the ODE solver can be adjusted by setting the step size $\mathrm{d}t=1/K$. This increases the sampling speed, but FMPE is not designed to optimize few-step sampling performance. We confirm this in our experiments with a rectified flow that encourages straight-path solutions (Liu et al., 2022).
@@ 2.4 Neural Posterior Score Estimation
Another approach to simulation-based inference with free-form networks lies in neural posterior score estimation (NPSE). This direction uses conditional score-based diffusion models (Song & Ermon, 2019; Song et al., 2021) to learn the posterior distribution. Sharrock et al. (2022) propose sequential NPSE, which trains a score-based diffusion model in a sequential (non-amortized) manner. With a slightly different focus, (Geffner et al., 2022) factorize the posterior distribution and learn scores of the (diffused) posterior for subsets (down to a single observation) of a larger data set. Subsequently, the information from the subsets is aggregated by combining the learned scores to approximate the posterior distribution of the entire data set.
Crucially, both methods rely on the basic formulation of score-based diffusion models: They gradually diffuse the target distribution through a diffusion process $\{\boldsymbol{\theta}_{t}\}_{t\in[0,T]}$ via a stochastic differential equation,
@@
with drift coefficient $\mu$, diffusion coefficient $\sigma$, time $t\in[0,T]$, and Brownian motion $\{\mathbf{w}_{t}\}_{t\in[0,T]}$. At each time $t$, the current (diffused) distribution of $\boldsymbol{\theta}$ conditional on $\mathbf{x}$ is denoted as $p_{t}(\boldsymbol{\theta}\,|\,\mathbf{x})$. Crucially, the distribution at $t=0$ equals the target posterior distribution, $p_{0}(\boldsymbol{\theta}\,|\,\mathbf{x})\equiv p(\boldsymbol{\theta}\,|\,\mathbf{x})$, and the distribution at $t=T$ is pure noise (see below). Song et al. (2021) prove that there exists an ordinary differential equation (“Probability Flow ODE”) whose solution trajectories at time $t$ are distributed according to $p_{t}(\boldsymbol{\theta}\,|\,\mathbf{x})$:
@@
where $\nabla\log p_{t}(\boldsymbol{\theta}_{t}\,|\,\mathbf{x})$ is the score function of $p_{t}(\boldsymbol{\theta}\,|\,\mathbf{x})$. This differential equation is usually designed to yield a spherical Gaussian noise distribution $p_{T}(\boldsymbol{\theta}\,|\,\mathbf{x})\approx\mathcal{N}(\mathbf{0},T^{2}\mathbf{I})$ after the diffusion process. Since we do not have access to the target posterior $p(\boldsymbol{\theta}\,|\,\mathbf{x})$, score-based diffusion models train a time-dependent score network $s_{\boldsymbol{\phi}}(\boldsymbol{\theta}_{t},t,\mathbf{x})\approx\nabla\log p_{t}(\boldsymbol{\theta}_{t}\,|\,\mathbf{x})$ via score matching and insert it into Eq. 4. Setting $\mu(\boldsymbol{\theta}_{t},t;\,\mathbf{x})=0$ and $\sigma_{t}=\sqrt{2t}$, the estimate of the Probability Flow ODE becomes $\mathrm{d}\boldsymbol{\theta}_{t}=-ts_{\boldsymbol{\phi}}(\boldsymbol{\theta}_{t},\mathbf{x},t)\mathrm{d}t$ (Karras et al., 2022). Finally, we can generate a random draw from the noise distribution $\boldsymbol{\theta}_{T}\sim\mathcal{N}(\mathbf{0},T^{2}\mathbf{I})$ and solve the Probability Flow ODE backwards for a trajectory $\{\boldsymbol{\theta}_{t}\}_{t\in[T,0]}$. The end of the trajectory $\boldsymbol{\theta}_{0}$ represents a draw from the approximate posterior $p_{0}(\boldsymbol{\theta}_{0}\,|\,\mathbf{x})\approx p(\boldsymbol{\theta}\,|\,\mathbf{x})$.
Numerical stability
The solver is usually stopped at a fixed small positive number $t=\varepsilon$ to prevent numerical instabilities (Song et al., 2023), so we use $\boldsymbol{\theta}_{\varepsilon}$ to denote the draw from the approximate posterior. For simplicity, we will also refer to $\boldsymbol{\theta}_{\varepsilon}$ as the trajectory’s origin.

### v3

@@ Abstract
Simulation-based inference (SBI) is constantly in search of more expressive and efficient algorithms to accurately infer the parameters of complex simulation models. In line with this goal, we present consistency models for posterior estimation (CMPE), a new conditional sampler for SBI that inherits the advantages of recent unconstrained architectures and overcomes their sampling inefficiency at inference time. CMPE essentially distills a continuous probability flow and enables rapid few-shot inference with an unconstrained architecture that can be flexibly tailored to the structure of the estimation problem. We provide hyperparameters and default architectures that support consistency training over a wide range of different dimensions, including low-dimensional ones which are important in SBI workflows but were previously difficult to tackle even with unconditional consistency models. Our empirical evaluation demonstrates that CMPE not only outperforms current state-of-the-art algorithms on hard low-dimensional benchmarks, but also achieves competitive performance with much faster sampling speed on two realistic estimation problems with high data and/or parameter dimensions.
@@ 1 Introduction
Simulation-based inference (SBI) comprises a family of computational methods for modeling the hidden properties of complex systems by means of simulation [1, 2]. In recent years, deep learning – and generative models in particular – have proven indispensable for scaling up SBI to challenging inverse problems [3, 4, 5, 6, 7]. Recently, multiple streams of neural SBI research have been capitalizing on the rapid progress in generative modeling of unstructured data by re-purposing existing generative architectures into general inverse problem solvers for applications in the sciences [8, 9, 10, 11]. In line with the above trend, we contribute a new contender to the growing suite of SBI methods for amortized Bayesian inference.
@@
Within the class of deep generative models, score-based diffusion models [12, 13, 14, 15] and flow matching algorithms [16, 17] have recently attracted significant attention due to their sublime performance as realistic generators. Diffusion models and flow matching are strikingly flexible but require a relatively expensive multi-step sampling phase to denoise samples [12]. To address this shortcoming, Song et al. [18] proposed consistency models (CMs), which are trained to perform few-step generation by design and have since emerged as a standalone generative model family. Thus, CMs appear attractive as a backbone architecture for SBI workflows due to their unique combination of unconstrained architectures and fast sampling speed – a conceptual advantage over the current state-of-the-art in neural SBI.
However, paradoxically, CMs may struggle with low-dimensional estimation problems, which are less important for image generation but crucial in typical SBI applications: These SBI applications are often characterized by relatively low-dimensional parameter spaces, even though the conditioning vectors (i.e., observables) might be high-dimensional [e.g., 3, 19, 20, 21]. Additionally, the quality of conditional CMs as few-step Bayesian samplers has not yet been explored empirically (e.g., in terms of probabilistic calibration and precision), even though this is crucial for their application in science and engineering. Lastly, while CMs for image generation are trained on enormous amounts of data, training data are typically scarce in SBI applications. In our empirical evaluations, we demonstrate that CMs are competitive with state-of-the-art SBI algorithms in low-data regimes using our adjusted settings (see Appendix A for details).
In this paper, we propose conditional CMs for SBI along with generalizable hyperparameter settings that achieve competitive performance with much faster sampling speed on both challenging low-dimensional benchmarks and high-dimensional applied problems. Our main contributions are:
@@
    We showcase the fundamental advantages of consistency models for amortized simulation-based inference: expressive unconstrained architectures and fast inference;
@@
    We demonstrate that CMPE outperforms NPE (normalizing flows) on three benchmark experiments (see Figure 1 above), is competitive with FMPE (flow matching) on Bayesian image denoising, and surpasses both NPE and FMPE on a challenging scientific simulator of tumor spheroid growth.
@@ 2 Preliminaries and related work
This section recaps simulation-based inference (SBI) via normalizing flows, flow matching, and score-based diffusion models. Readers familiar with these topics can safely fast-forward to Section 3.
2.1 Notation
In the following, the neural network training relies on a synthetic training set $\{(\boldsymbol{\theta}^{(m)},\mathbf{x}^{(i)})\}_{m=1}^{M}$, which consists of parameter-data tuples. Each superscript $m$ marks one training example, namely, a tuple of a latent parameter vector and an observable data set which was generated from the parameter vector. The total simulation budget for training the neural networks follows as $M$. We summarize the $D$-dimensional latent parameter vector of the simulator as $\boldsymbol{\theta}\equiv(\theta_{1},\ldots,\theta_{D})$. Further, each data set $\mathbf{x}\equiv\{\mathbf{x}_{n}\}_{n=1}^{N}$ is a matrix whose rows consist of $N$ vector-valued observations. Accordingly, one parameter vector $\boldsymbol{\theta}^{(m)}$ yields one data set $\mathbf{x}^{(m)}$ from the simulator (see Section 2.2). In Bayesian inference, we aim to infer the unknown parameters $\boldsymbol{\theta}$ of the mechanistic simulator, which are not to be confused with the trainable neural network weights $\boldsymbol{\phi}$.
@@ 2.2 Simulation-based inference (SBI)
Computer simulations play a fundamental role across countless scientific disciplines, ranging from physics to biology, and from climate science to economics [2]. Simulators $g(\cdot,\cdot)$ generate observables $\mathbf{x}\in\mathcal{X}$ as a function of unknown parameters $\boldsymbol{\theta}\in\Theta$ and latent program states $\boldsymbol{\xi}\in\Xi$ [1]:
$\mathbf{x}=g(\boldsymbol{\theta},\boldsymbol{\xi})\quad\text{with}\;\boldsymbol{\theta}\sim p(\boldsymbol{\theta}),\;\boldsymbol{\xi}\sim p(\boldsymbol{\xi}\,|\,\boldsymbol{\theta})$ (1)
The forward problem in Equation 1 is typically well-understood through scientific theories instantiated as generative models. The inverse problem, however, is much harder, and forms the crux of Bayesian inference: reduce uncertainty about the unknowns $\boldsymbol{\theta}$ based on observables $\mathbf{x}$ through the posterior distribution $p(\boldsymbol{\theta}\,|\,\mathbf{x})\propto p(\boldsymbol{\theta})\,p(\mathbf{x}\,|\,\boldsymbol{\theta})$. The key property of neural SBI and early approximate Bayesian computation [ABC; 22, 23] methods is that they approach the inverse problem by sampling (i.e., simulating) from the forward model. Thus, the forward model acts as an implicit statistical model [24] that enables proper Bayesian inference in the limit of infinite simulations. In contrast, likelihood-based methods (e.g., MCMC) depend on the explicit evaluation of the likelihood $p(\mathbf{x}\,|\,\boldsymbol{\theta})$.
We can classify neural SBI methods into sequential [e.g., 25, 26, 27, 28, 29, 10] and amortized [e.g., 30, 5, 31, 32, 33, 11, 9, 34, 35, 10] methods. Sequential inference algorithms iteratively refine the prior $p(\boldsymbol{\theta})$ to generate simulations in the vicinity of the target observation. Thus, they are not amortized, as each new observed data set requires a potentially costly re-training of the neural approximator tailored to the particular target observation. However, most sequential methods can be turned amortized by training the neural approximator to generalize over the entire prior predictive space of the model. This allows us to query the approximator with any new data set during inference. In fact, amortization can be performed across any component of the model, including multiple data sets [5] and contextual factors, such as the number of observations in a data set [31], heterogeneous data sources [36], or even different probabilistic models [37, 38]. Our current work is situated in the amortized setting.
@@ 2.3 Normalizing flows for neural posterior estimation
Neural posterior estimation (NPE) methods for SBI have traditionally relied on conditional discrete normalizing flows for learning a conditional neural density estimator $p_{\boldsymbol{\phi}}(\boldsymbol{\theta}\,|\,\mathbf{x})$ from simulated pairs $(\boldsymbol{\theta},\mathbf{x})$ of parameters and data [39, 8]. Normalizing flows can learn closed-form approximate densities via maximum likelihood training:
@@
Two prominent normalizing flow architectures in SBI are affine coupling flows [ACF; 40] and neural spline flows [NSF; 41]. A major disadvantage of these architectures is their strict bijectivity requirement: it restricts the design space of the neural networks to invertible functions with cheap Jacobian calculations, hence the desire for more flexible architectures. In the following, we will briefly recap recent works that have pioneered the use of multi-step, unconstrained architectures for SBI, inspired by the success of such models in a variety of generative tasks [15, 13].
@@ 2.4 Flow matching for posterior estimation
Wildberger et al. [11] applied flow matching techniques [16, 17] in SBI, an approach abbreviated as flow matching posterior estimation (FMPE). FMPE is based on optimal transport [42, 43], where the mapping between base and target distribution is parameterized by a continuous process driven by a vector field $\mu$ on the sample space for each time step $t\in[0,1]$. Here, the distribution at $t=1$ could be a unit Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I})$ and the distribution at $t=0$ is the target posterior. The FMPE loss replaces the maximum likelihood term in the NPE objective (Eq. 2) with a conditional flow matching objective,
@@
where $\mu_{\boldsymbol{\phi}}$ denotes the conditional vector field parameterized by a unconstrained neural network with trainable parameters $\boldsymbol{\phi}$, and $u_{t}$ denotes a marginal vector field, which can be as simple as $u_{t}(\boldsymbol{\theta}_{t}\,|\,\boldsymbol{\theta})=\boldsymbol{\theta}_{1}-\boldsymbol{\theta}$ for all $t$ [17]. We obtain posterior draws $\boldsymbol{\theta}_{0}\sim p(\boldsymbol{\theta}\,|\,\mathbf{x})$ by solving $\mathrm{d}\boldsymbol{\theta}_{t}=-\mu(\boldsymbol{\theta}_{t},t;\mathbf{x})$ in reverse on $t\in[0,1]$, starting with noise samples $\boldsymbol{\theta}_{1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. We can use any off-the-shelf ODE solver for transforming noise $\boldsymbol{\theta}_{1}$ into a draw $\boldsymbol{\theta}_{0}$ from the approximate posterior. In principle, the number of steps $K$ in the ODE solver can be adjusted by setting the step size $\mathrm{d}t=1/K$. Decreasing the number of steps increases the sampling speed, but FMPE is not designed to optimize few-step sampling performance. We confirm this in our experiments with a rectified flow that encourages straight-path solutions [17].
@@ 2.5 Neural posterior score estimation
Another approach to simulation-based inference with unconstrained networks lies in neural posterior score estimation [NPSE; 10], which can either feature single-round inference (amortized) or multiple sequential inference rounds on a particular data set (non-amortized). This stream of research uses conditional score-based diffusion models [44, 12] to learn the posterior distribution. With a slightly different focus, [9] factorize the posterior distribution and learn scores of the diffused posterior for subsets (down to a single observation) of a larger data set. Subsequently, the information from the subsets is aggregated by combining the learned scores to approximate the posterior distribution of the entire data set. Crucially, both methods rely on the basic formulation of score-based diffusion models: They gradually diffuse the target distribution according to the following diffusion process,
@@
with drift coefficient $\mu$, diffusion coefficient $\sigma$, time $t\in[0,T]$, and Brownian motion $\{\mathbf{w}_{t}\}_{t\in[0,T]}$. At each time $t$, the current (diffused) distribution of $\boldsymbol{\theta}$ conditional on $\mathbf{x}$ is denoted as $p_{t}(\boldsymbol{\theta}\,|\,\mathbf{x})$. Crucially, the distribution at $t=0$ equals the target posterior distribution, $p_{0}(\boldsymbol{\theta}\,|\,\mathbf{x})\equiv p(\boldsymbol{\theta}\,|\,\mathbf{x})$, and the distribution at $t=T$ is set to be Gaussian noise (see below). Song et al. [12] prove that there exists an ordinary differential equation (“Probability Flow ODE”) whose solution trajectories at time $t$ are distributed according to $p_{t}(\boldsymbol{\theta}\,|\,\mathbf{x})$,
@@
where $\nabla\log p_{t}(\boldsymbol{\theta}_{t}\,|\,\mathbf{x})$ is the score function of $p_{t}(\boldsymbol{\theta}\,|\,\mathbf{x})$. This differential equation is usually designed to yield a spherical Gaussian noise distribution $p_{T}(\boldsymbol{\theta}\,|\,\mathbf{x})\approx\mathcal{N}(\mathbf{0},T^{2}\mathbf{I})$ after the diffusion process. Since we do not have access to the target posterior $p(\boldsymbol{\theta}\,|\,\mathbf{x})$, score-based diffusion models train a time-dependent score network $s_{\boldsymbol{\phi}}(\boldsymbol{\theta}_{t},t,\mathbf{x})\approx\nabla\log p_{t}(\boldsymbol{\theta}_{t}\,|\,\mathbf{x})$ via score matching and insert it into Eq. 5. Setting $\mu(\boldsymbol{\theta}_{t},t;\,\mathbf{x})=0$ and $\sigma_{t}=\sqrt{2t}$, the estimate of the Probability Flow ODE becomes $\mathrm{d}\boldsymbol{\theta}_{t}=-ts_{\boldsymbol{\phi}}(\boldsymbol{\theta}_{t},\mathbf{x},t)\mathrm{d}t$ [45]. Finally, we can generate a random draw from the noise distribution $\boldsymbol{\theta}_{T}\sim\mathcal{N}(\mathbf{0},T^{2}\mathbf{I})$ and solve the Probability Flow ODE backwards for a trajectory $\{\boldsymbol{\theta}_{t}\}_{t\in[T,0]}$. The end of the trajectory $\boldsymbol{\theta}_{0}$ represents a draw from the approximate posterior $p_{0}(\boldsymbol{\theta}_{0}\,|\,\mathbf{x})\approx p(\boldsymbol{\theta}\,|\,\mathbf{x})$.
Numerical stability. The solver is usually stopped at a fixed small positive number $t=\varepsilon$ to prevent numerical instabilities [18], so we use $\boldsymbol{\theta}_{\varepsilon}$ to denote the draw from the approximate posterior. For simplicity, we will also refer to $\boldsymbol{\theta}_{\varepsilon}$ as the trajectory’s origin.